x/5 - x+3/x = -3/x
some help to solve this pleeeeeease

Answer by stanbon(75887)

x/5 - (x+3)/x = -3/x
------------
the least common multiply of the denominators is 5x.
Multiply thru by 5x to get:
x^2 - 5(x+3) = -3*5
x^2-5x-15=-15
x^2-5x=0
x(x-5)=0
x=0 or x=5
But x cannot be zero as it is in the denominator or
the original problem.
So, x=5 is the only answer.
Cheers,
Stan H.



==============
Cheers,
Stan H.